# Retaining by Doing: The Role of On-Policy Data in Mitigating Forgetting

### v1

@@ 4.1 On-Policy Data is the Primary Contributor for Mitigating Forgetting
Table 1: REINFORCE also enjoys a lesser degree of forgetting as GRPO. Gain and drop comparison of SFT, REINFORCE, and GRPO. The SFT and GRPO results are taken from Figure 2.
@@ 4.2 Approximately On-Policy Data Can Suffice for Mitigating Forgetting
We have identified the usage of on-policy data as the main contributor for the robustness of RL to forgetting. However, generating on-policy data at each step entails a non-negligible compute overhead. We therefore explore the degree of “on-policyness” required to enjoy the benefit of reduced forgetting. Figure 2 showed that Self-SFT, which generates data only from the initial policy, suffers from severe forgetting. On the other hand, RL, which generates data at every step and thus represents the most on-policy end of the spectrum, is robust to forgetting. We now test whether Iterative-SFT, an approximately on-policy approach that iteratively trains on data generated at the start of each epoch (Zelikman et al., 2022; Dong et al., 2023; Xiong et al., 2025), can suffice for mitigating forgetting.
@@ 5 Related Work
Catastrophic forgetting. Catastrophic forgetting has been studied since early research on connectionist models (McCloskey & Cohen, 1989). Early studies mostly focused on preventing parameters from drastically changing (Kirkpatrick et al., 2017; Li & Hoiem, 2018; Lopez-Paz & Ranzato, 2017). In the context of LM post-training, patterns of catastrophic forgetting differ due to the massive data used during pre-training (Luo et al., 2023; Shi et al., 2024; Wu et al., 2024). While LMs typically do not drastically forget all pre-trained knowledge, post-training LMs often leads to degradation in performance, which has been called “alignment tax” (Bai et al., 2022; Ouyang et al., 2022). Prior work found more severe forgetting when considering contrasting domains such as instruction following and safety (Qi et al., 2023; He et al., 2024). Though, Lee et al. (2024); Chen et al. (2024) suggest that forgotten behaviors or abilities can be revived with little re-training. Kotha et al. (2024) posited that forgetting happens when the LM infers a wrong mode from the mixture of distributions to perform the task, and a carefully selected prompt can recover forgetting. Our work draws inspiration from the mixture of distributions perspective to establish intuition.
LM post-training. Post-training techniques such as SFT and RL are widely used for endowing pre-trained LMs with desired behaviors or enhancements in target domains. SFT relies on the ground truth responses provided in the datasets. RL allows the model to generate its own responses and only provide a reward at the end, be it from a parameterized reward model (Schulman et al., 2017) or a verifiable reward module (Shao et al., 2024; Lambert et al., 2024). Recent studies have shown that SFT and RL exhibit distinct characteristics. Chu et al. (2025) showed that RL is able to generalize to unseen distributions while SFT mostly memorizes. (Wang et al., 2025) observed that RL can benefit from even training on a single example without severe overfitting. (Mukherjee et al., 2025) reported that RL naturally fine-tunes a smaller network compared to SFT. A common thread connecting these observations is that the parameter update during RL training is more local and targeted. Other methods such as RAFT (Dong et al., 2023) and STaR (Zelikman et al., 2022) performs SFT in several rounds and can be viewed as on-policy offline RL. This paper complements these studies and provide a forgetting-centric view on the difference between SFT and RL.
@@ 7 Acknowledegment
We thank the members of Princeton Language Intelligence Group for providing comments on the manuscript. We also thank Kevin Lu from Thinking Machines Lab for the valuable feedback. This research is supported by the National Science Foundation (IIS-2211779) and is also supported in part by Schmidt Sciences. NR is supported in part by the Zuckerman STEM Leadership Program.
@@ A.5 KL Divergence From Initial Policy and Forgetting Correlate Moderately
Shenfeld et al. (2025) have concurrently found that RL exhibits less forgetting than SFT, attributing RL’s robustness to an implicit regularization of RL toward policies with low KL divergence from the initial policy $\pi_{\theta_{0}}$. In particular, Shenfeld et al. (2025) empirically identify $\mathrm{KL}[\,\pi_{\theta_{0}}\,||\,\pi_{\theta}\,]$ as an indicator for the extent of forgetting, mostly through an extensive evaluation on synthetic tasks. We explore the connection between $\mathrm{KL}[\,\pi_{\theta_{0}}\,||\,\pi_{\theta}\,]$ and forgetting in the setup of §2, where the KL divergence is estimated based on $100$ examples from the evaluation set. Table 2 shows that, in accordance with the hypothesis of Shenfeld et al. (2025), GRPO exhibits both smaller KL divergences and smaller drops in non-target tasks performance compared to the SFT variants. Furthermore, the Pearson correlation between KL divergence and non-target tasks drop across all models, methods, and datasets is $0.52$. Yet, when comparing Self-SFT and SFT, the relation between KL divergence and forgetting is less monotonic—a larger KL does not necessarily imply a higher degree of forgetting. This indicates that the relationship between KL divergence and forgetting may be more subtle than Shenfeld et al. (2025) suggest.

### v2

@@ 4.1 On-Policy Data is the Primary Contributor for Mitigating Forgetting
Table 1: The advantage estimate of GRPO is not responsible for its robustness to forgetting. Comparison of SFT, REINFORCE, and GRPO. The SFT and GRPO results are taken from Figure 2.
@@ 4.2 Approximately On-Policy Data Can Suffice for Mitigating Forgetting
We identified the usage of on-policy data as the main contributor for the robustness of RL to forgetting. However, generating on-policy data at each step entails a non-negligible compute overhead. We therefore explore the degree of “on-policyness” required to enjoy the benefit of reduced forgetting. Figure 2 showed that Self-SFT, which generates data only from the initial policy, suffers from severe forgetting. On the other hand, RL, which generates data at every step and thus represents the most on-policy end of the spectrum, is robust to forgetting. We now test whether Iterative-SFT, an approximately on-policy approach that iteratively trains on data generated at the start of each epoch (Zelikman et al., 2022; Dong et al., 2023; Xiong et al., 2025), can suffice for mitigating forgetting.
@@ 5 Related Work
Catastrophic forgetting. Catastrophic forgetting has been studied since early research on connectionist models (McCloskey & Cohen, 1989). Initial efforts to mitigate forgetting focused on preventing parameters from drastically changing (Kirkpatrick et al., 2017; Li & Hoiem, 2018; Lopez-Paz & Ranzato, 2017). In the context of LM post-training, patterns of catastrophic forgetting differ due to the massive data used during pre-training (Luo et al., 2023; Shi et al., 2024; Wu et al., 2024). While LMs typically do not drastically forget all pre-trained knowledge, post-training LMs often leads to degradation in performance, which has been called “alignment tax” (Bai et al., 2022; Ouyang et al., 2022). Prior work found more severe forgetting when considering contrasting domains such as instruction following and safety (Qi et al., 2023; He et al., 2024). Though, Lee et al. (2024); Chen et al. (2024) suggest that forgotten behaviors or abilities can be revived with little re-training. Kotha et al. (2024) posited that forgetting happens when the LM infers a wrong mode from the mixture of distributions to perform the task, and a carefully selected prompt can recover forgetting. Our work draws inspiration from the mixture-of-distributions perspective to establish intuition.
LM post-training. Post-training methods such as SFT and RL are widely used for endowing pre-trained LMs with desired behaviors or enhancing performance on target tasks. SFT relies on the ground truth demonstrations. By contrast, RL generates responses from the model and only provides a reward signal, be it from a parameterized reward model (Schulman et al., 2017) or verifiable rewards (Shao et al., 2024; Lambert et al., 2024). Recent studies have shown that SFT and RL exhibit distinct characteristics. Razin et al. (2024; 2025) identified that the optimization speed of RL strongly depends on reward variance. Chu et al. (2025) showed that RL is able to generalize to unseen distributions while SFT can be prone to memorization. Wang et al. (2025) observed that RL can benefit from even training on a single example without severe overfitting. Lastly, Mukherjee et al. (2025) reported that RL fine-tunes a smaller network compared to SFT. A common thread connecting these results is that the parameter update during RL training is more local and targeted. Other methods such as RAFT (Dong et al., 2023) and STaR (Zelikman et al., 2022) perform SFT in several rounds and can be viewed as approximately on-policy RL. This paper complements these studies and provides a forgetting-centric view on the difference between SFT and RL.
@@ 7 Acknowledegment
We thank the members of Princeton Language Intelligence Group for providing comments on the manuscript. We also thank Kevin Lu from Thinking Machines Lab for the valuable feedback. This research is supported by the National Science Foundation (IIS-2211779), Cisco Research, and is also supported in part by Schmidt Sciences. NR is supported in part by the Zuckerman STEM Leadership Program.
@@ A.5 KL Divergence From Initial Policy and Forgetting Correlate Moderately
Shenfeld et al. (2025) have concurrently found that RL exhibits less forgetting than SFT, attributing RL’s robustness to an implicit regularization of RL toward policies with low KL divergence from the initial policy $\pi_{\theta_{0}}$. In particular, Shenfeld et al. (2025) empirically identify $\mathrm{KL}[\,\pi_{\theta_{0}}\,||\,\pi_{\theta}\,]$ as an indicator for the extent of forgetting, mostly through an extensive evaluation on synthetic tasks. We explore the connection between $\mathrm{KL}[\,\pi_{\theta_{0}}\,||\,\pi_{\theta}\,]$ and forgetting in the setup of §2, where the KL divergence is estimated based on $100$ examples from the evaluation set. Table 2 shows that, in accordance with the hypothesis of Shenfeld et al. (2025), GRPO exhibits both smaller KL divergences and smaller drops in non-target tasks performance compared to the SFT variants. Furthermore, the Pearson correlation between KL divergence and non-target tasks drop across all models, methods, and datasets is $0.52$. Yet, when comparing Self-SFT and SFT, the relation between KL divergence and forgetting is less monotonic—a larger KL does not necessarily imply a higher degree of forgetting. This indicates that the relationship between KL divergence and forgetting is still not fully understood.